OpenStudy (anonymous):

Hello I could really use help with this definite integral!! Thank You! ʃ (secx(tanx + secx) dx (0 to pi/4)

OpenStudy (anonymous):

This is the question

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}(\sec(x)\tan(x) + \sec^2(x))dx = \sec(x) + \tan(x)[0\rightarrow \pi/4]\]\[= \sec(0) + \tan(0) - \sec(\pi /4) - \tan(\pi / 4) = 1 + 0 - \sqrt{2} - 1 = - \sqrt{2}\]

OpenStudy (anonymous):

did it backwards. answer is + sqrt(2)

OpenStudy (anonymous):

should be\[\sec(\pi/4)+\tan(\pi/4)−\sec(0)−\tan(0)= \sqrt2\]

OpenStudy (anonymous):

thank you!!